OpenStudy (anonymous):

√5:(2x^2+1)-2 =0 how can I solve this. Step by step please.

OpenStudy (tkhunny):

You'll have to provide a better expression. What's that colon doing in there?

OpenStudy (anonymous):

I have a logarithmic equation that I really need help with ASAP. This is the problem log(x^2+4) - log(x+2)= 2 + log(x-2)

OpenStudy (tkhunny):

Have you considered changing 2 into log(100)?

OpenStudy (anonymous):

can you help me break it down into step by step?

OpenStudy (anonymous):

this is the answer: \[2/3\sqrt{101/11}\] but I just do not know how it get to this.

OpenStudy (tkhunny):

I did. I gave you an interesting first step. Use log rules and demonstrate the next.

OpenStudy (anonymous):

aight. Thank you.